A salesperson's weekly paycheck is 25% less yjan her co-worker's paycheck. The two paychecks total $750. find the amount of each paycheck. (Round your answer to 2 decimal places.)
Answer by checkley77(12569)

x+.75x=750
1.75x=750
x=750/1.75
x=428.57 for the larger pay check.
Proof:
428.57+.75*428.57=750
428.57+321.43=750
750=750